HC2H3O2 (aq) + NaHCO3 (aq) → NaC2H3O2 (aq) + H2O (l) + CO2 (g)
1. Calculate the amount of vinegar (acetic acid (HC2H3O2) 4.25%) and 3 M sodium bicarbonate (NaHCO3) solution required to produce enough carbon dioxide (CO2) to fill a 2.5 L container at 2 atm of pressure and 122 °C.
2. Calculate the density of the sample of carbon dioxide (CO2) collected from the previous
questions. Plot a graph (density and temperature) showing the density changes as the
temperature decreases to 101, 78, 53, 33 and 19°C.
3. Calculate the enthalpy of the reaction. Use Wolfram Alpha for the standard enthalpies of
formation.
1
Expert's answer
2019-05-17T02:42:59-0400

1. PV=nRT

n=PV/RT = 2 x 2.5 / 0.082 x 395=0.1543 mol

2. d= PM / RT = 2 x 44 / 0.082 x 395 = 2.71688793 g/ml

3. 2.2 kj/mol